3x-1 over x-8 subtract -4 < 0 and the sign has the line under it

3x^2=4x-11

Found 2 solutions by richwmiller, solver91311:
Answer by richwmiller(17219)


((3x-1)/(x-8))-4<0
Which sign has what line under it.
Do you mean the < less than is less than or equal?
You need to learn how to write math problems.
Math is language with its own words and grammar.
You can't just ignore it ans expect anyone to understand you
What does that have to do with 3^2=4x-11
So do you have a question?
Reminder:
You are limited to one problem per submission.
No more than 4 submissions daily
No similar problems.

Answer by solver91311(24713)



I think you mean:



In which case you should have written:

((3x - 1)/(x - 8)) - 4 <= 0

And then there would have been no confusion.

On the other hand, you might have meant



because you said "subtract -4", in which case you should have written:

((3x - 1)/(x - 8)) - (-4) <= 0

Presuming the first rendering is correct and presuming, though you didn't say so, that you want to solve for , proceed as follows:

Add 4 to both sides of the inequality:



Multiply both sides by



Add and 1 to both sides:



Multiply both sides by -1 and reverse the sense of the inequality:




John